# MgFNO: Multi-grid Architecture Fourier Neural Operator for Parametric Partial Differential Equations

### v1

@@ 3.2 Compensation Methods of Low-Frequency Tendency
In order to address the issues caused by slow convergence at high frequencies and speed up the training of FNO, several methods have been proposed. Jagtap and Ameya D [27] used activation functions $\sigma(\mu ax)$ instead of $\sigma(\mu x)$. Here, $\mu$ is a fixed scaling factor, and $a$ is a trainable parameter for all neurons. Cai and Wei [28] introduced PhaseDNN, which converts the high-frequency components of data to the low-frequency spectrum for learning, and then converts the learned frequency components back to the original high-frequency spectrum. Subsequently, the multiscale DNN method (MscaleDNN) was proposed [22, 29], which can be understood as Fourier expansion only in the radial direction. The frequency space conversion can be achieved through scaling, which is equivalent to inverse scaling in spatial space. In this article, we utilize the conversion of coarse and fine grids and MscaleDNN to address the slow convergence issue of high-frequency errors in FNO. Next, let us review the ideas and structure of MscaleDNN in regular DNNs.
@@
A common MscaleDNN structure is shown in Figure 7. We use neural networks successively in physical space to learn the stretched sub-functions $f_{i}^{(scale)}(\alpha_{i}x)$. Furthermore, it is mentioned in [Liu_2020] that when constructing MscaleDNN, activation functions with compact support should be selected. In MgFNO, we select the following function $\phi(x)$ as our activation function.
@@
Figure 7: : A common MscaleDNN structure
On the other hand, looking back at the MG method we mentioned in Section 2, it is worth noting that iterative methods (2.1) tend to eliminate high-frequency error components rather than low-frequency ones. Therefore, we may employ a straightforward approach to remove these low-frequency error components. As it is shown in figure 8, for MG method on a coarser mesh, the low frequency components transform into high frequency ones. Because if there is an error on one side of the mesh and a different error on the other side of the mesh, once we coarsen the cells and then bring them next to each other, the variation between those two positions is now a high spatial frequency. Then we can remove high-frequency components with a few iterations of iterative methods. Due to the opposite convergence tendency of DNNs and iterative methods in the frequency domain, we have decided to apply a multi-grid architecture similar to the MG method to FNO, but this time we need to transfer the coarse grid to the fine grid. We hope that through the design of this multi-grid architecture, FNO can overcome the high-frequency issues and accelerate the model training process.
@@ 5 Numerical Implementation and Results
Data Set and Result. Firstly, we set the viscosity coefficient $\nu=0.1$. Then, for initial condition $u_{0}(x)$, we set $u_{0}\sim\mu$ where $\mu=\mathcal{N}(0,625(-\Delta+25I)^{-2})$. Finally, we solve the equation using a split step method where the heat equation part is solved exactly in Fourier space then the non-linear part is advanced, again in Fourier space, using a very fine forward Euler method. We solve on a spatial mesh with resolution $81922^{13}=$ 8192, where lower resolution data are downsampled from.

### v2

@@ 3.2 Compensation Methods of Low-Frequency Tendency
To address the issues caused by slow convergence at high frequencies and speed up the training of FNO, several methods have been proposed. Jagtap and Ameya D [27] used activation functions $\sigma(\mu ax)$ instead of $\sigma(\mu x)$. Here, $\mu$ is a fixed scaling factor, and $a$ is a trainable parameter for all neurons. Cai and Wei [28] introduced PhaseDNN, which converts the high-frequency components of data to the low-frequency spectrum for learning, and then converts the learned frequency components back to the original high-frequency spectrum. Subsequently, the multiscale DNN method (MscaleDNN) was proposed [22, 29], which can be understood as Fourier expansion only in the radial direction. The frequency space conversion can be achieved through scaling, which is equivalent to inverse scaling in spatial space. In this article, we utilize the conversion of coarse and fine grids and MscaleDNN to address the slow convergence issue of high-frequency errors in FNO. Next, let us review the ideas and structure of MscaleDNN in regular DNNs.
@@
A common MscaleDNN structure is shown in Figure 7. We use neural networks successively in physical space to learn the stretched sub-functions $f_{i}^{(scale)}(\alpha_{i}x)$. Furthermore, it is mentioned in [22] that when constructing MscaleDNN, activation functions with compact support should be selected. In MgFNO, we select the following function $\phi(x)$ as our activation function.
@@
Figure 7: : A common MscaleDNN structure. Reprinted from Liu et al. [22]
On the other hand, looking back at the MG method we mentioned in Section 2, it is worth noting that iterative methods (2.1) tend to eliminate high-frequency error components rather than low-frequency ones. Therefore, we may employ a straightforward approach to remove these low-frequency error components. As it is shown in figure 8, for the MG method on a coarser mesh, the low-frequency components transform into high-frequency ones. If there is an error on one side of the mesh and a different error on the other side of the mesh, once we coarsen the cells and then bring them next to each other, the variation between those two positions is now a high spatial frequency. Then we can remove high-frequency components with a few iterations of iterative methods. Due to the opposite convergence tendency of DNNs and iterative methods in the frequency domain, we have decided to apply a multi-grid architecture similar to the MG method to FNO, but this time we need to transfer the coarse grid to the fine grid. We hope that through the design of this multi-grid architecture, FNO can overcome the high-frequency issues and accelerate the model training process.
@@ 5 Numerical Implementation and Results
Data Set and Result. Firstly, we set the viscosity coefficient $\nu=0.1$. Then, for initial condition $u_{0}(x)$, we set $u_{0}\sim\mu$ where $\mu=\mathcal{N}(0,625(-\Delta+25I)^{-2})$. Finally, we solve the equation using a split step method where the heat equation part is solved exactly in Fourier space then the non-linear part is advanced, again in Fourier space, using a very fine forward Euler method. We solve on a spatial mesh with resolution $81922^{13}=$ 8192, where lower resolution data are down-sampled from.